OpenStudy (anonymous):

tyree sold tee shirts .he sold six tee shirts which was 30% of his invintory how many t shirts are in his invintory?

OpenStudy (mathstudent55):

You can use a proportion. 6 shirts are 30% of something. What number is 100% of the same something. 6 is to 30 as x is to 100 6:30 = x:100 \(\dfrac{6}{30} = \dfrac{x}{100} \) Solve for x.

OpenStudy (anonymous):

wait i still dont get it

OpenStudy (mathstudent55):

Have you learned proportions?

OpenStudy (anonymous):

kinda i need help on the question can you explain better i am doing homework thats why i ask

OpenStudy (mathstudent55):

Ok. Here is another way of looking at it. In math, the word "of" usually means "times." Also, percent, or %, means divide by 100.

OpenStudy (mathstudent55):

The question your problem asks is: 6 is 30% of a number. Find the number.

OpenStudy (anonymous):

AND HOW DO IMPLY I DO THAT AIR HEAD

OpenStudy (anonymous):

YOU ARE NOT LISTINING

OpenStudy (mathstudent55):

We use the math vocabulary above to translate the question into an equation. BTW, the word "is" usually means =. Also, for "a number" which means an unknown number, we use a variable which is usually a letter, very commonly x. English: 6 is 30% of a number. Math: 6 = 30/100 * x

OpenStudy (mathstudent55):

Now we have a math equation that we can solve to find the number you need. \(6 = \dfrac{30}{100}x\) Solve for x.

OpenStudy (anonymous):

BYE AND BUY A MATH BOOK GO BACK TO SCHOOL

OpenStudy (anonymous):

LOOK I AM NOT HAVING A nighnteen hour convesation message when your sober

OpenStudy (anonymous):

good by

OpenStudy (mathstudent55):

\(6 = \dfrac{30}{100}x\) \(6 = \dfrac{3}{10}x\) Multiply both sides by \(\dfrac{10}{3} \): \(\dfrac{10}{3} \times 6 = \dfrac{10}{3} \times \dfrac{3}{10}x\) \(\dfrac{60}{3} = x\) \(x = 20\) The answer is: His inventory had 20 shirts. He sold 6 shirts, so his inventory has 14 shirts now.

OpenStudy (mathstudent55):

Are you happy now? It was way less than 19 hours.

OpenStudy (anonymous):

sorry that was my older sister being mean i needed a drink real quik we share rooms no privacy what so ever

OpenStudy (anonymous):

wait what is sober